# When does bottom-up beat top-down in hierarchical community detection?

### v2

@@ 3.1 Hierarchical Stochastic Block Model
The hierarchical stochastic block model (HSBM) is a class of random graphs whose nodes are partitioned into latent hierarchical communities. Before defining this model formally, let us introduce some notations. Each node $u$ of a rooted binary tree $\mathcal{T}$ is represented by a binary string as follows. The root is indexed by the empty string $\emptyset$. Each non-root node $u$ of the tree is labeled by the unique binary string $u=u_{1}u_{2}\cdots u_{q}\cdots$ that records the path from the root (i.e., $u_{q}=1$ if step $q$ of the path is along the right branch of the split and $u_{q}=0$ otherwise). The depth of node $u$ is denoted by $|u|$ and coincides with its distance from the root [^2] By convention, $|\emptyset|=0$.. Finally, with this parametrization, the lowest common ancestor $\mathrm{lca}(u,v)$ of two nodes $u,v\in\mathcal{T}$ is simply the longest common prefix of the binary strings $u$ and $v$.
@@
We suppose that the probability $p_{ab}$ of having a link between two nodes belonging to the primitive communities $C_{a}$ and $C_{b}$ depends only on the lowest common ancestor of $a$ and $b$. Hence, we denote the probability $p_{ab}$ by $p(\mathrm{lca}(a,b))$. In an assortative setting, we have $p(u)<p(v)$ if $v$ is a child of $u$. Finally, we need to rule out settings in which the difference $p(v)-p(u)$ decreases too fast with respect to $p(u)$. To avoid this flatness of $p$ issue, we always assume that lim $\liminf\frac{\left|p(v)-p(u)\right|}{\max\{p(u),p(v)\}}>0$ for any child $v$ of a parent $u$. We now define an assortative HSBM.
@@
Let $N$ be a positive integer, $\mathcal{T}$ a rooted binary tree with $|\mathcal{L}_{\mathcal{T}}|=K$ leaves, $\pi=(\pi_{a})_{a\in\mathcal{L}_{\mathcal{T}}}$ a probability vector and a function $p\colon\mathcal{T}\to[0,1]$ such that for $u,v\in\mathcal{T}$ we have $p(u)<p(v)$ (assortativity) if $v$ is a child of $u$. A hierarchical stochastic block model (HSBM) is a graph $G=(V,E)$ such that $V=[N]$ and
@@
An important particular case of HSBM is the binary tree SBM (BTSBM), in which the tree $\mathcal{T}$ is full and balanced, and the probability of a link between two nodes in clusters $C_{a}$ and $C_{b}$ depends only on the depth of $\mathrm{lca}(a,b)$, i.e., $p(\mathrm{lca}(a,b))=p_{|\mathrm{lca}(a,b)|}$ for all $a,b\in\mathcal{L}_{\mathcal{T}}$. In particular, the number of communities of a BTSBM is $K=2^{d_{\mathcal{T}}}$, and assortativity implies that $p_{0}<p_{1}<\cdots<p_{d_{\mathcal{T}}}$. We illustrate an HSBM and a BTSBM in Figure 1.
(a) Hierarchical stochastic block model.
(b) Binary tree SBM of depth 3.
Figure 1: Examples of an HSBM (a) and a BTSBM (b), with the binary string representation of each node. The link probabilities are $p(u)$ for the HSBM and $p_{|u|}$ for the BTSBM. The grey-colored rectangles represent the super-communities.
3.2 Tree Recovery from the Bottom Clusters
@@
where $B\,\Delta\,C=\left(B\cup C\right)\backslash\left(B\cap C\right)$ denotes the symmetric difference between two sets $B,C\subset V$. The minimum is taken over the set $\mathcal{S}_{[K]}$ of all permutations of $[K]$, because we recover only the bottom communities up to a global permutation of the community labels. We study sequences of networks indexed by the number of nodes $N$ and for which the interaction probabilities might depend on $N$. An estimator $\widehat{\mathcal{C}}$ of $\mathcal{C}$ achieves almost exact recovery if $\mathbb{P}\left(\mathrm{loss}\left(\mathcal{C},\widehat{\mathcal{C}}\right)=o(N)\right)=1-o(1)$ and exact recovery if $\mathbb{P}\left(\mathrm{loss}\left(\mathcal{C},\widehat{\mathcal{C}}\right)=0\right)=1-o(1)$.
The number of edges between two communities $C_{a}$ and $C_{b}$ is binomially distributed with mean $|C_{a}|\cdot|C_{b}|\cdot p_{ab}$. Therefore, if $p_{ab}=\omega(N^{-2})$, this binomial random variable is concentrated around its mean and we have $\rho(C_{a},C_{b})=(1+o(1))p_{ab}$. This suggests that, once the primitive communities are almost exactly recovered, the average-linkage procedure successfully recovers the tree from its leaves if the edge probability between different communities is $\omega\left(N^{-2}\right)$. The proof is more involved because (i) we compute $\rho(\widehat{C}_{a},\widehat{C}_{b})$ and not $\rho(C_{a},C_{b})$ and (ii) the estimator $\widehat{\mathcal{C}}$ is correlated with the graph structure. In particular, to establish that $\rho(\widehat{C}_{a},\widehat{C}_{b})=(1+o(1))p_{ab}$, we impose $p_{ab}=\omega(N^{-1})$.
@@
Consider an assortative HSBM whose latent binary tree $\mathcal{T}$ has $K$ leaves. Suppose that $K$ is constant, $\min_{a\in\mathcal{L}_{\mathcal{T}}}\pi_{a}>0$, and $p(t)=a(t)\delta_{N}$ where $(a(t))_{t\in\mathcal{T}}$ are positive constants (independent of $N$) and $N\delta_{N}=\omega(1)$. Let $\widehat{\mathcal{C}}$ be an estimator of $\mathcal{C}$, possibly correlated with the graph edges and such that $|\widehat{\mathcal{C}}|=|\mathcal{C}|$. If $\widehat{\mathcal{C}}$ is almost exact, then the average-linkage procedure correctly recovers $\mathcal{T}$ (starting from $\widehat{\mathcal{C}}$).
@@
We prove Theorem 1 in Appendix A.1. The assumption of the existence of an almost exact estimator $\widehat{\mathcal{C}}$ of $\mathcal{C}$ is not limiting here. Indeed, under the assumptions of Theorem 1, we can obtain such an estimator by the spectral algorithm of YP (16) or by the agnostic-degree-profiling algorithm of AS15b .
3.3 Robustness of the Linkage Procedure
In this section, we study the robustness of the linkage procedure to classification errors made by the initial clustering, when the number of errors scale as $\Theta(N)$. For simplicity, we restrict the study to BTSBMs.
Suppose that each node has a probability $\eta$ to be misclustered and that misclusterings occur uniformly across all cluster pairs. More precisely, the probability that a node $i\in[N]$ belonging to block $a$ is misclustered to a different block $b\neq a$ depends only on the depth of the lowest common ancestor between $a$ and $b$. We denote this probability by $\zeta(|\mathrm{lca}(a,b)|)$. Finally, let $\widetilde{\mathcal{C}}(\mathcal{C},\zeta)$ be the bottom clusters obtained by this procedure [^3] Note that $\widetilde{\mathcal{C}}(\mathcal{C},\zeta)$ depends on $\zeta$ and $\mathcal{C}$ but is independent of the graph edges. This is unrealistic because the linkage procedure starts from the predicted bottom clusters found by applying a clustering algorithm on the graph (see Algorithm 1). From a theoretical perspective, we can split the original graph $G=(V,E)$ into two subgraphs $G_{1}=(V,E_{1})$ and $G_{2}=(V,E_{2})$ such that $E_{1}\cup E_{2}=E$. An edge $(u,v)\in E$ is in $E_{1}$ with probability $\gamma$, otherwise it is in $E_{2}$. We apply the flat clustering algorithms on $G_{1}$ to obtain $\widehat{\mathcal{C}}^{1}$ and run the linkage procedure on $G_{2}$; $\widehat{\mathcal{C}}^{1}$ is independent of $E_{2}$..
A natural assumption is that nodes are more likely to be misclustered in blocks close to their true block; this is verified if $\zeta$ is monotonically non-decreasing.
@@ 3.3 Robustness of the Linkage Procedure
Let $G$ be a BTSBM whose latent binary tree $\mathcal{T}$ has depth $d$ and $K=2^{d}$ leaves. Suppose that $\pi_{a}=1/K$ for all $a\in\mathcal{L}_{\mathcal{T}}$ and that $\min\limits_{u\in\mathcal{T}}p(u)=\omega(N^{-2})$. Assume that $\zeta$ is non-decreasing and verifies $(\zeta(d-1)+\zeta(d-2))/2<1-\eta$. Let $\widetilde{\mathcal{C}}=\widetilde{\mathcal{C}}(\mathcal{C},\zeta)$ be a clustering obtained as described at the beginning of Section 3.3. Then the average-linkage procedure correctly recovers $\mathcal{T}$ (starting from $\widetilde{\mathcal{C}}$).
Example 1 shows that if (i) the misclustered nodes are assigned to a cluster uniformly at random among all $K$ clusters and if (ii) the proportion $1-\eta$ of correctly clustered nodes is strictly larger than $1/K$, then average-linkage recovers the tree. As $1/K$ is the expected proportion of correctly clustered nodes resulting from a random guess, Condition (ii) is weak.
@@
Let $\zeta_{\mathrm{uniform}}(h)=\eta/(K-1)$ for $10\leq h\leq d-$ 1. If $1-\eta>1/K$, then the average-linkage recovers $\mathcal{T}$ starting from $\widetilde{\mathcal{C}}_{\mathrm{uniform}}=\widetilde{\mathcal{C}}(\mathcal{C},\zeta_{\mathrm{uniform}})$.
$\zeta_{\mathrm{adversarial}}(h)\ =\ \begin{cases}\frac{\eta}{2^{d-1}}&\text{ if }h=0,\\ 0&\text{ if }h\in\{1,\cdots,d-1\}.\end{cases}$ (3.2)
Proposition 1.
$(i)\ \bar{p}_{1}<p_{d-1}\qquad\text{ or }\qquad(ii)\ \bar{p}_{1}\geq p_{d-1}\text{ and }\eta<\eta_{-}.$
@@ 4 Exact Recovery at Intermediate Depths
We now focus on the exact recovery of communities at intermediate depths within the hierarchy and establish tight information-theoretic conditions determining the feasibility of achieving exact recovery at a specific depth. It is intuitively expected that recovering super-communities of higher depths should be comparatively easier than the recovery of lower depths. Therefore, we expect scenarios where the exact recovery of the primitive communities might be unattainable, but where the exact recovery of super-communities at intermediate depths remains achievable. To provide context, we initially recapitulate key findings on exact recovery in non-hierarchical stochastic block models (SBMs) in Section 4.1. We then present our main results in Section 4.2, which gives the precise conditions for exact recovery at intermediate depths.

### v3

@@ 3.1 Hierarchical Stochastic Block Model
The hierarchical stochastic block model (HSBM) is a class of random graphs whose nodes are partitioned into latent hierarchical communities. Before defining this model formally, let us introduce some notations. Each node $u$ of a rooted binary tree $\mathcal{T}$ is represented by a binary string as follows. The root is indexed by the empty string $\emptyset$. Each non-root node $u$ of the tree is labeled by the unique binary string $u=u_{1}u_{2}\cdots u_{q}\cdots$ that records the path from the root (i.e., $u_{q}=1$ if step $q$ of the path is along the right branch of the split and $u_{q}=0$ otherwise). The depth of node $u$ is denoted by $|u|$ and coincides with its distance from the root. Finally, with this parametrization, the lowest common ancestor $\mathrm{lca}(u,v)$ of two nodes $u,v\in\mathcal{T}$ is the longest common prefix of the binary strings $u$ and $v$. This is also the common ancestor of $u$ and $v$ with the largest depth.
@@
We suppose that the probability $p_{ab}$ of having a link between two nodes belonging to the primitive communities $C_{a}$ and $C_{b}$ depends only on the lowest common ancestor of $a$ and $b$. Hence, we denote the probability $p_{ab}$ by $p(\mathrm{lca}(a,b))$. In an assortative setting, we have $p(u)<p(v)$ if $v$ is a child of $u$.
@@
Let $N$ be a positive integer, $\mathcal{T}$ a rooted binary tree with $|\mathcal{L}_{\mathcal{T}}|=K$ leaves, and $\pi=(\pi_{a})_{a\in\mathcal{L}_{\mathcal{T}}}$ a probability vector. Let $p\colon\mathcal{T}\to[0,1]$ be an increasing function, meaning that for any $u,v\in\mathcal{T}$, if $v$ is a child of $u$ we have $p(u)<p(v)$. A hierarchical stochastic block model (HSBM) is a graph $G=(V,E)$ such that $V=[N]$ and
@@
An important particular case of HSBM is the binary tree SBM (BTSBM), in which the tree $\mathcal{T}$ is full and balanced, and the probability of a link between two nodes in clusters $C_{a}$ and $C_{b}$ depends only on the depth of $\mathrm{lca}(a,b)$, i.e., $p(\mathrm{lca}(a,b))=p_{|\mathrm{lca}(a,b)|}$ for all $a,b\in\mathcal{L}_{\mathcal{T}}$. In particular, the number of communities of a BTSBM is $K=2^{d_{\mathcal{T}}}$, and assortativity implies that $p_{0}<p_{1}<\cdots<p_{d_{\mathcal{T}}}$. We illustrate an HSBM and a BTSBM in Figure 1. Observe that the BTSBM defines the same model as in li2022hierarchical, while the HSBM is a natural extension where the tree is still binary but not necessarily full and complete.
(a) Hierarchical stochastic block model
(b) Binary tree SBM of depth 3
Figure 1: Examples of (a) an HSBM and (b) a BTSBM, with the binary string representation of each node. The link probabilities are $p(u)$ for the HSBM and $p_{|u|}$ for the BTSBM. The grey-colored rectangles represent the super-communities.
In our results, we use the following assumptions.
Assumption 1 (Fixed hierarchy).
$\mathcal{T}$ is a rooted binary tree with $K$ leaves and is independent of $N$. The probability vector $\pi\in(0,1)^{K}$ is also independent of $N$ and satisfies $\min_{a\in K}\pi_{a}>0$.
Assumption 2 (Asymptotic scaling).
The edge connection probabilities $p$ can be written as $p(t)=a(t)\delta_{N}$, where $(a(t))_{t\in\mathcal{T}}$ are positive constants (independent of $N$).
Under Assumption 2, the average degree is $\Theta(N\delta_{N})$, and we call $\delta_{N}$ the sparsity factor.
3.2 Tree Recovery with Growing Average Degree
@@
where $B\,\Delta\,C=\left(B\cup C\right)\backslash\left(B\cap C\right)$ denotes the symmetric difference between two sets $B,C\subset V$. The minimum is taken over the symmetric group $\mathcal{S}_{[K]}$ of all permutations of $[K]$, because we recover the bottom communities only up to a global permutation of the community labels. We study sequences of networks indexed by the number of nodes $N$ and for which the interaction probabilities might depend on $N$. An estimator $\widehat{\mathcal{C}}$ of $\mathcal{C}$ achieves exact recovery if $\mathrm{loss}\left(\mathcal{C},\widehat{\mathcal{C}}\right)\xrightarrow{\rm{a.s.}}0$, almost exact recovery if $N^{-1}\mathrm{loss}\left(\mathcal{C},\widehat{\mathcal{C}}\right)\xrightarrow{p}0$ and weak recovery if $\mathbb{P}(N^{-1}\mathrm{loss}\left(\mathcal{C},\widehat{\mathcal{C}}\right)<\|\pi\|^{2}_{2}-\varepsilon)\to 0$ for all $\varepsilon>0$. [^2] The interpretation is as follows. An exact estimator makes no mistakes, while an almost exact estimator makes $o(N)$ mistakes. A weak estimator makes $O(N)$ mistakes but still outperforms a naive random guessing, which ignores the graph structure and classifies a node $i$ into community $a$ with probability $\pi_{a}$, independently of other nodes.
The number of edges between two communities $C_{a}$ and $C_{b}$ is binomially distributed with mean $|C_{a}|\cdot|C_{b}|\cdot p_{ab}$. Therefore, if $p_{ab}=\omega(N^{-2})$, this binomial random variable is concentrated around its mean and we have $\rho(C_{a},C_{b})=(1+o(1))p_{ab}$. This suggests that, once the primitive communities are almost exactly recovered, the average-linkage procedure successfully recovers the tree from its leaves if the edge probability between different communities is $\omega\left(N^{-2}\right)$. The proof is more involved because (i) we compute $\rho(\widehat{C}_{a},\widehat{C}_{b})$ and not $\rho(C_{a},C_{b})$, and (ii) the estimator $\widehat{\mathcal{C}}$ is correlated with the graph structure. In particular, to establish that $\rho(\widehat{C}_{a},\widehat{C}_{b})=(1+o(1))p_{ab}$, we impose $p_{ab}=\omega(N^{-1})$.
@@
Consider an assortative HSBM. Suppose that Assumptions 1 and 2 hold, with $N\delta_{N}=\omega(1)$. Let $\widehat{\mathcal{C}}$ be an estimator of $\mathcal{C}$, possibly correlated with the graph edges and such that $|\widehat{\mathcal{C}}|=|\mathcal{C}|$. If $\widehat{\mathcal{C}}$ is almost exact, then Algorithm 1 recovers $\mathcal{T}$ (starting from $\widehat{\mathcal{C}}$).
@@
We prove Theorem 1 in Appendix A.1. The assumption of the existence of an almost exact estimator $\widehat{\mathcal{C}}$ of $\mathcal{C}$ is not limiting here. Indeed, under the assumptions of Theorem 1, we can obtain such an estimator by the spectral algorithm of yun2016optimal or by the agnostic-degree-profiling algorithm of abbe2015recovering. Because the average degree is of the order $\Theta(N\delta_{N})$, Theorem 1 ensures that Algorithm 1 recovers the hierarchy when the average degree grows unbounded. [^3] When $N\rho_{N}=\omega(1)$, the graph may be disconnected; however, only $o(N)$ nodes are excluded from the largest connected component, making almost exact recovery possible (but not exact recovery).
3.3 Tree Recovery with Bounded Average Degree
For simplicity, throughout this section, we assume that $\mathcal{T}$ is a full and balanced tree, and the edge probabilities satisfy $p(t)=a(t)\delta_{N}$ where $(a(t))_{t\in\mathcal{T}}$ are positive constants (independent of $N$), with $N\delta_{N}=\Theta(1)$. This bounded average degree regime is more challenging, as no algorithm can achieve almost exact recovery. [^4] When $\delta_{N}=\Theta(1/N)$ and the average degree exceeds 1, there exists a single giant component containing $\Theta(N)$ vertices, while all other components are of size $o(N)$. Importantly, this giant component includes vertices from all $K$ communities; thus, there is no scenario in which two communities are entirely disconnected. Therefore, even in this extremely sparse regime, it remains possible to infer the community structure and estimate inter-community edge densities. We have clarified this point in the revised text to aid reader understanding. However, when $N\delta_{N}\geq C$ where $C$ is a quantity depending only on $a(t)$ but independent of $N$, some algorithms achieve weak recovery abbe2017community. This leads to the following question: can average-linkage recover the tree, but starting with a weak estimator $\widehat{\mathcal{C}}$ of $\mathcal{C}$ instead of an almost exact estimator?
To answer this question, we must handle the fact that a weak estimator misclassifies $\Theta(N)$ nodes, as opposed to $o(N)$ for an almost exact estimator. Recall from the discussion before Theorem 1 that the correlation between $\widehat{\mathcal{C}}$ and the graph $G$ makes it challenging to study the concentration of $\rho(\widehat{C}_{a},\widehat{C}_{b})$. As a result, we can no longer guarantee that $\rho(\widehat{C}_{a},\widehat{C}_{b})=(1+o(1))p_{ab}$ when $\widehat{\mathcal{C}}$ is only a weak estimator. To mitigate this issue, we split $G$ into two graphs $G^{1}$ and $G^{2}$ and proceed in two steps. We first apply the flat-clustering algorithm on $G^{1}$ to obtain $\widehat{\mathcal{C}}^{1}$, and next estimate the edge density by
$\rho_{2}(\widehat{C}_{a}^{1},\widehat{C}_{b}^{1})\ =\ \frac{\sum_{\begin{subarray}{c}i\in\widehat{C}_{a}^{1},j\in\widehat{C}_{b}^{1}\end{subarray}}A_{ij}^{2}}{\left|\widehat{C}_{a}^{1}\right|\cdot\left|\widehat{C}_{b}^{1}\right|},$ (3.2)
where $A^{2}$ is the adjacency matrix of $G^{2}$. We call this two-step technique graph-splitting.
Definition 2 (Graph-splitting).
The graph-splitting of a graph $G=(V,E)$ with split-probability $\gamma$ generates two random graphs $G_{1}=(V,E_{1})$ and $G_{2}=(V,E_{2})$ having the same node set as $G$. The graph $G_{1}$ is formed by independently sampling each edge of $G$ with probability $\gamma$. The graph $G_{2}$ consists of the remaining edges that were not includesd in $G_{1}$,i.e., $E_{2}=E\backslash E_{1}$.
When $G$ is an HSBM with edge connection probabilities $p_{ab}$, the subgraphs $G_{1}$ and $G_{2}$ are also HSBMs with the same community structure as $G$, but with respective edge connection probabilities $\gamma p_{ab}$ and $(1-\gamma)p_{ab}$. As a result, graph-splitting provides an estimator $\widehat{\mathcal{C}}^{1}$ that is now independent of $G_{2}$, which solves the issue of correlation between $\widehat{\mathcal{C}}$ and $G$. However, this estimator still misclassifies $\Theta(N)$ nodes, and we make the following simplifying assumption regarding the misclassified nodes.
Assumption 3.
Assume that there exists a weakly consistent estimator $\widehat{\mathcal{C}}^{1}=(\widehat{C}_{1},\cdots,\widehat{C}_{K})$, that is an estimate of $\mathcal{C}=(C_{1},\cdots,C_{K})$, and let $O_{ab}=C_{a}\cap\widehat{\mathcal{C}}^{1}_{b}$ be the nodes in cluster $C_{a}$ but assigned to cluster $\widehat{\mathcal{C}}^{1}_{b}$. We suppose that
$\frac{O_{ab}}{N/K}\ =\ (1+o(1))\,\zeta(|\mathrm{lca}(a,b)|),$
where $\zeta$ is a non-decreasing function of the depth satisfying $\zeta(d-1)+\zeta(d-2)<2\zeta(d)$.
Because all communities of a BTSBM have expected size $N/K$, the quantity $\frac{O_{ab}}{N/K}$ represents the fraction of vertices in cluster $a$ that are clustered in cluster $b$. Hence, the assumption that $\zeta$ is non-decreasing implies that a node is more likely to be clustered into a cluster closer to its true cluster rather than one further away in the tree. Because $a=b$ if and only if $|\mathrm{lca}(a,b)|=d$, the quantity $\zeta(d)$ is the probability that a node is correctly clustered. In contrast, $\zeta(h)$ for $0\leq h\leq d-1$ is the probability that a node from cluster $a$ is misclustered into cluster $b$ whose least common ancestor with $a$ is at depth $h$. The assumption $\zeta(d-1)+\zeta(d-2)<2\zeta(d)$ ensures that the proportion of misclustered nodes is not too large, and is automatically verified if $\zeta$ is strictly increasing. The following theorem establishes that Algorithm 1 successfully recovers the hierarchy, even when starting with an estimator that makes $\Theta(N)$ mistakes.
@@ 3.3 Tree Recovery with Bounded Average Degree
Let $G$ be a BTSBM. Suppose that Assumptions 1 and 2 hold and $N\delta_{N}=\Theta(1)$. Let $G_{1}$ and $G_{2}$ be the two graphs obtained from $G$ by graph-splitting with any $\gamma$ satisfying $\gamma=1-o(1)$ and $1-\gamma=\omega(1/N)$. Let $\widehat{\mathcal{C}}^{1}$ be a clustering obtained from $G^{1}$ and satisfying Assumption 3. Then Algorithm 1 correctly recovers $\mathcal{T}$ (using the edge density defined in Equation (3.2)).
The assumption $1-\gamma=\omega(1/N)$ ensures that the edge density of $G_{2}$ is $\omega(N^{-2})$, and hence that $G_{2}$ contains enough edges so that $\rho_{2}(\widehat{C}_{a}^{1},\widehat{C}_{b}^{1})$ defined in Equation (3.2) concentrates around its expectation. The condition $\gamma=1-o(1)$ ensures that if weak recovery is possible from $G$, then it also remains possible from $G_{1}$: in other words, graph-splitting does not result in a loss of information.
@@
Denote by $\eta\geq 0$ the proportion of misclustered nodes and suppose that $\zeta(0)=\cdots=\zeta(d-1)$. This imposes that $\zeta(d)=1-\eta$ and $\zeta(h)=\eta/(K-1)$ for $0\leq h\leq d-1$. The condition $\zeta(d-1)+\zeta(d-2)<2\zeta(d)$ in Assumption 3 is equivalent to $\eta<1-1/K$.
@@ 4 Exact Recovery at Intermediate Depths
We now focus on the exact recovery of communities at intermediate depths within the hierarchy and establish tight information-theoretic conditions determining the feasibility of achieving exact recovery at a specific depth. It is intuitively expected that recovering super-communities of smaller depths should be comparatively easier than the recovery at larger depths. Therefore, we expect scenarios where the exact recovery of the primitive communities might be unattainable, but where the exact recovery of super-communities at intermediate depths remains achievable. To provide context, we initially recapitulate key findings on exact recovery in non-hierarchical stochastic block models (SBMs) in Section 4.1. We then present our main results in Section 4.2, which gives the precise conditions for exact recovery at intermediate depths. Although the results of this section do not specifically involve hierarchical algorithms, we discuss in Section 5.2 implications of exact recovery at intermediate depths for the theoretical guarantees of top-down algorithms.